OpenStudy (anonymous):

Its a continuity question. I'm having problem of finding the 'a'. Question, is

OpenStudy (anonymous):

\[\tan ax / \tan bx = 4\]given b =4

OpenStudy (anonymous):

\[\tan ax \over \tan bx \]

OpenStudy (anonymous):

@Rohangrr oh yea, the question is to find 'a'.

OpenStudy (anonymous):

Are you sure there are no limits ??

OpenStudy (anonymous):

yes there is, limit is x < 0 do you want me to give the exact question? because this one i already simplify it

OpenStudy (anonymous):

PLease do give the exact question

OpenStudy (anonymous):

tan(ax )= tan(bx) implies that ax=kPi +bx maybe this helps...

OpenStudy (anonymous):

ok got it :)

OpenStudy (anonymous):

a=16

OpenStudy (anonymous):

Using this property \[\lim_{x \rightarrow 0} tanx /x =1 \]

OpenStudy (anonymous):

yes a=16 @shivam_bhalla , so the after step is?

OpenStudy (anonymous):

Sorry the answer is a= 16 The question is limit x->0 (tan ax) / (ax) * (bx) / (tan bx) * (ax)/(bx) = 4 now limit a* x->0 (tan ax) / (ax) = 1 limit b* x->0 (tan bx) / (bx) = 1 Therefore 1*1* ax/bx=4 a/b=4 a= 16

OpenStudy (anonymous):

Sorry for deleting posts so many times :P